Multiply by whole numbers (lot of)

Multiplying by a whole number always means finding that number of the quantity.

1. What fraction is the same length as 3 lots of 2/9?

(It should be clear that it is 2/3)

2. What fraction is the same length as 4 lots of 1/12?

3. What fraction is the same as 6 lots of 2/15?


1
Expert's answer
2021-06-29T08:01:04-0400

1. What fraction is the same length as 3 lots of 2/9?


It follows that required fraction is equal to "3\cdot\frac{2}{9}=\frac{6}{9}=\frac{2}{3}."


2. What fraction is the same length as 4 lots of 1/12?


It follows that required fraction is equal to "4\cdot\frac{1}{12}=\frac{4}{12}=\frac{1}{3}."


3. What fraction is the same as 6 lots of 2/15?


It follows that required fraction is equal to "6\cdot\frac{2}{15}=\frac{12}{15}=\frac{4}{5}."